# Simultaneously Constraining the Neutron Star Equation of State and Mass Distribution through Multimessenger Observations and Nuclear Benchmarks

### v2

@@ VI.2 Impact of Including Less Reliable Pulsar Mass Measurements on EOS Constraints
Figure 9 compares the 90% credible regions of the mass-radius posterior distributions for two cases: (i) using only highly reliable pulsar mass measurements, combined with astrophysical and nuclear constraints, and (ii) including both reliable and less reliable mass measurements along with the same astrophysical and nuclear inputs. We observe that the inclusion of less reliable pulsar mass measurements leads to significantly tighter constraints on the EOS.
Specifically, the inferred maximum mass of a neutron star increases slightly and is better constrained with the inclusion of less reliable data. The maximum mass is determined to be $2.25^{+0.13}_{-0.10}\,M_{\odot}$, a much narrower credible interval compared to the results obtained using only reliable mass measurements. This improvement highlights the potential of leveraging less reliable measurements, despite their inherent uncertainties, to refine EOS constraints.
The tighter constraint is likely due to the additional data points spanning a broader range of masses and radii, which effectively reduce the parameter space for viable EOS models. This result demonstrates that, while reliability is critical for robust inference, carefully incorporating less reliable data can provide complementary insights and improve overall constraints.
@@
Figure 9: Comparison of mass-radius posterior distributions using only reliable pulsar mass measurements (blue) and including less reliable measurements (orange). Both cases include additional astrophysical observations and nuclear physics inputs. The shaded regions represent the 90% credible intervals.

### v3

@@ VI.2 Impact of Including Less Reliable Pulsar Mass Measurements on EOS Constraints
Figure 10 compares the 90% credible regions of the mass-radius posterior distributions for two cases: (i) using only highly reliable pulsar mass measurements, combined with astrophysical and nuclear constraints, and (ii) including both reliable and less reliable mass measurements along with the same astrophysical and nuclear inputs. We observe that the inclusion of less reliable pulsar mass measurements leads to significantly tighter constraints on the EOS. Specifically, the inferred maximum mass of a neutron star increases slightly and is better constrained with the inclusion of less reliable data. The maximum mass is determined to be $2.25^{+0.13}_{-0.10}\,M_{\odot}$, a much narrower credible interval compared to the results obtained using only reliable mass measurements. This improvement highlights the potential of leveraging less reliable measurements, despite their inherent uncertainties, to refine EOS constraints. The tighter constraint is likely due to the additional data points spanning a broader range of masses and radii, which effectively reduce the parameter space for viable EOS models. This result demonstrates that, while reliability is critical for robust inference, carefully incorporating less reliable data can provide complementary insights and improve overall constraints.
@@
Figure 10: Comparison of mass-radius posterior distributions using only reliable pulsar mass measurements (orange) and including less reliable measurements (blue). Both cases include additional astrophysical observations and nuclear physics inputs. The shaded regions represent the 90% credible intervals.